OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x2 − 8x − 42

OpenStudy (paxpolaris):

\[2x^2-8x-42\] each has 2 in common, so: \[=2(x^2-4x-21)\]

OpenStudy (paxpolaris):

\[-7+3=-4\]\[-7\cdot3=-21\]

OpenStudy (anonymous):

thank you:)

OpenStudy (paxpolaris):

\[=2(x−7)(x+3)\]